### v1

@@ 2.3 Impact on thermodynamic derivatives
Figure 3: Relative difference of the thermodynamic derivative $\nabla_{\rm ad}$ between the two EoS’, D-EoS and T-EOS, in log scale. The composition used is that of pure $\rm{}^{16}O$, $\overline{A}=16$ and $\overline{Z}=8$. The numbered zones represented are the same as in Figs. 1 and  2.
Between different EoSs the thermodynamic derivatives can change significantly. Specifically we focus on the derivative needed to compute the adiabatic gradient,
@@
This must be obtained with sufficient accuracy, since $\nabla_{\rm ad}$ is the fundamental variable determining whether a mass shell is convectively unstable. A region is deemed convective if the radiative gradient, $\nabla_{\rm rad}=d\ln T/\rm d\ln P$, is greater than $\nabla_{\rm ad}$.
@@ 3 Extension of the opacity coverage
The current implementation of opacities for the advanced phases in GENEC does not cover the full $(\rho,T)$ plane needed to evolve massive star models until core collapse. We shall use analytic extensions to estimate the opacities in these high energy and relativistic regimes.
@@
In the following we detail the extensions employed to compute both $\kappa_{\rm cond}$ and $\kappa_{\rm rad}$ in the high density and high temperature regime, which is reached as stellar models evolve to core-collapse.
@@ 3.1 Radiative opacity
For most of the $(\rho,T)$ plane the radiative opacities are evaluated from tabulated values (see Ekström et al. 2012, for further details), which take into account many physical processes such as bound-bound transitions, bound-free transitions, free-free transitions and electron scattering to name just a few. The coverage of these tables generally only extend to temperatures of $10^{9}$ K. At higher temperatures, the opacity is dominated by electron scattering, but the Klein-Nishina corrections to Compton scattering lower it (Weaver et al. 1978), and at even higher temperatures, the effects of electron-positron pairs must be included in the opacity (Woosley et al. 2002). In the high density regime, electron conduction is an important contribution to the opacity (see Itoh et al. 1985), and also electron degeneracy tends to reduce the opacity due to occupation of the electron phase space. To extend the computation of the radiative opacity to these regions, we apply the work of Poutanen (2017) to estimate the Rosseland mean opacity for electron scattering in the high-temperature regime. This opacity is expressed as,
$\kappa_{\rm es}=\frac{\sigma_{T}N_{e}}{\rho}\frac{1}{\Lambda}.$ (8)
where the parameter $\Lambda(T,\psi)$ is obtained from a fit in temperature and degeneracy parameter. The fit employed is as follows,
@@ 3.2 Conductive opacity
The conductive opacities due to electrons are computed using the formulation of Iben (1975), given in the Appendix of that paper. These were already employed in the previous version of GENEC (Ekström et al. 2012), however here the expressions have been modified to apply also in the final phases of stellar evolution. For this reason, we detail next the procedure used to compute $\kappa_{\rm cond}$. There are two main cases, the first being when the electrons are non-relativistic, roughly for $\rho<10^{6}\,\text{g\,cm}^{-3}$. In that case, a fit is developed based on the results of Hubbard & Lampe (1969), where they compute the thermal electron conductivity for mixtures of pure hydrogen, helium, and carbon. Thus, for densities smaller than $10^{6}\,\text{g\,cm}^{-3}$ the conductive opacity is written as
@@
$Z_{C}=\frac{1}{3}\sum_{i}Z_{i}^{2}X_{i}/A_{i},$ (15)
for all species beyond helium, ($T_{6}=T/10^{6}$). Here $\theta_{X}$, $\theta_{Y}$, $\theta_{C}$ and $f$ are fit parameters, dependent on thermodynamic quantities, as detailed in Iben (1975) to match the tabulated results of Hubbard & Lampe (1969). Note that the factor 1/3 in Eq. 15 is there to normalise the weight of $X_{\rm{}^{12}C}$ to 1 ($36^{2}/12=$ 3) because the computation for $\kappa_{\rm cond}$ in Hubbard & Lampe (1969) is only done for pure Carbon.
In the case of relativistic degenerate electrons, Iben (1975) proposes a fit for the results of Canuto (1970), which extend up to densities of $\rho\approx 10^{12}\,\text{g\,cm}^{-3}$. The approximation is,
$\kappa_{c}=6.35\times 10^{-8}\times\frac{T_{6}^{2}Z_{\beta}}{\epsilon_{F}^{2.5}G}\times\frac{1}{(1+\epsilon_{F})}\ \mathrm{cm^{2}\,g}^{-1}$ (16)
where,
$\epsilon_{F}$ $=\left[1+\left(\frac{\rho_{6}}{\mu_{e}}\right)^{2/3}\right]^{1/2}-1$
$Z_{\beta}$ $=\sum_{i}Z_{i}^{2}X_{i}/A_{i}$
$\log_{10}G$ $=[(0.873-0.298\log_{10}Z_{\alpha})$
$+(0.333-0.168\log_{10}Z_{\beta})M]\times\left[1-(1+\gamma)^{-0.85}\right]$
$Z_{\alpha}$ $=\sum_{i}Z_{i}^{2}X_{i}/A_{i}^{1/3}$
@@
$M$ $=\text{min}(1,0.5+\log_{10}\epsilon_{F}),$
with $\mu_{e}$ the electron chemical potential. These fits provide good coverage for large parts of the $(\rho,T)$ plane for generic compositions, and are generally accurate to the order of a few percent. In the latest version of the MESA code, the conductive opacities are computed from the tables of Cassisi et al. (2007) which have been extended to reach, $10^{10}\,\rm K$ in temperature, and $10^{11.5}\,\text{g\,cm}^{-3}$ in density (see Paxton et al. 2013). For our purposes, the fits provided by Iben (1975) estimate to sufficient accuracy the influence of electron conduction in the total opacity.
@@ 4 Implementation of an effective network
To reach the brink of core collapse, massive stars progressively neutronise their cores, reducing the electron fraction, $Y_{e}$, and with it the supporting degenerate electron pressure. Any reaction network accounting for the most advanced burning phases, must include (at least, approximately) the weak reactions that reduce the core electron fraction (mostly, electron captures onto elements of the iron group). In the previous work by Hirschi et al. (2004), GENEC models were computed post silicon core-burning and even post silicon shell burning, thus bringing them very close to collapse models. However, the network employed did not account for any decrease in $Y_{e}$. To extend the use of GENEC models as supernova progenitors, we have implemented electron captures in a reduced network to model the $Y_{e}$ decrease seen in stellar cores prior to collapse. Electron captures are especially important from silicon core burning to collapse (Woosley et al. 2002). Ideally, one would include a large set of reactions and isotopes to resolve with precision the $Y_{e}$ evolution. For instance, Rauscher et al. (2002) have employed a network of up to 2 20022002\,2002 200 isotopes, Farmer et al. (2016) using MESA computed the pre-supernova structure of models with various nuclear reaction networks of different sizes and Limongi & Chieffi (2018) used a network of roughly 300 isotopes.
In this work, we use a simple extension of the pre-existing $\alpha$-chain backbone network of GENEC. We approximate the effects of weak reactions that one would get in a larger network by taking one simple electron capture chain from $\rm{}^{56}Ni$ to $\rm{}^{56}Cr$. In this sense we create a similar network to the 21 species, detailed for example in Farmer et al. (2016), commonly refereed to in MESA literature (Paxton et al. 2011, 2013, 2015, 2018, 2019) as approx21_plus_co56. The exact network we use contains 25 species [^2] Note that in the Main Sequence and Helium core burning a larger network is used to capture CNO cycles correctly, see Ekström et al. (2012) for details. From Carbon burning onwards, we use the 25 species network described here. and is presented in Fig. 4. Hereafter, we will refer to this reduced network as GeValNet25. Compared with the approx21_plus_co56 network, we additionally include explicitly the iron group elements $\rm{}^{53}Fe$, $\rm{}^{55}Fe$, $\rm{}^{55}Co$, and $\rm{}^{57}Co$ to model the core neutronisation within the framework of GENEC.
@@
Note that the last electron capture in this chain contains a double step, first from $\rm{}^{56}Fe$ to $\rm{}^{56}Mn$, and then to $\rm{}^{56}Cr$. This electron chain mimics the effect of a much larger network containing many electron reactions (cf., e.g., Aufderheide et al. 1994; Langanke & Martínez-Pinedo 2000). The rates used here were calibrated in a way that we reach similar core electron fractions as previous studies (Weaver et al. 1978; Woosley et al. 2002; Heger & Woosley 2010; Chieffi et al. 1998; Farmer et al. 2016; Paxton et al. 2018). It should be noted that beta reactions are not explicitly included in our calculations. Hence, there is no true balance between weak processes, implying that once the chain of electron captures begins, it unavoidably brings the core towards a composition dominated by the last isotope of the chain. In other words, should the core not collapse, in a finite time all the core would be composed of $\rm{}^{56}Cr$. Thus, the electron capture rates have been suitably adapted to become effective weak rates that account for the state close to beta equilibrium in the core. As pointed out in Heger et al. (2001b), due to the quasi beta equilibrium in the core, the values of the beta decay rates are immaterial, as long as they facilitate the equilibrium. Indeed, even variations by factors 2 to 5 in the beta-decay rates do not significantly change the nuclear evolution. Note that the choice of using an electron capture chain along $A=56$ until $\rm{}^{56}Cr$ is not by chance. In studies with larger networks and more species, the final $Y_{e}$ value commonly falls around $0.43-0.45$ (Woosley et al. 2002). If the simplified electron capture chain stopped at $\rm{}^{56}_{26}Fe$, then the lowest $Y_{e}$ attainable would be $Y_{e}^{\rm min}=26/56\approx 0.46$ (when all the matter in a core mass shell would be converted to $\rm{}^{56}_{26}Fe$), too high to match the results of larger and more complete networks. However, including $\rm{}^{56}_{24}Cr$ in the chain allows one to evolve to $Y_{e}^{\rm min}=24/56\approx 0.42$. Thus, a core containing a mixture of $\rm{}^{56}Fe$ and $\rm{}^{56}Cr$ can reach the electron fraction values generally found with sophisticated studies, even without explicitly accounting for the balancing effect of beta-decays. We cautiously point out that, with a reduced network, we may not provide a complete forecast of the nuclear composition at the brink of collapse.
@@
Figure 4: Network used in all GENEC models from core C-burning onwards. Squares filled with a plus sign denote isotopes explicitly evolved in the nuclear reaction network. Squares filled with a blue dot are isotopes implicitly included in the $\alpha$-chain. By this we mean that along the $\alpha$-chain are included not only the ($\alpha$, $\gamma$) reaction but also the ($\alpha$, $p$)($p$, $\gamma$) reactions without explicitly evolving the intermediate isotope abundance. With such a network, we get a good estimation of the energy generation rate (see Timmes et al. 2000, for details). Additionally, compared to the previous GENEC network, we have the approximate electron captures from $\rm{}^{56}Ni$ to $\rm{}^{56}Cr$ to mimic pre-supernova $Y_{e}$ evolution. This network is very similar to that of the 22 isotopes network of Farmer et al. (2016), note that in Fig. 1 of that paper only 21 isotopes are flagged in the 22 isotope network, but also $\rm{}^{56}Co$ is added to the network making 22 in total. Compared with that network, we track explicitly extra iron group isotopes, namely, $\rm{}^{57}Co$ along with $\rm{}^{53}Fe$ and $\rm{}^{55}Fe$, thus, in total we track 25 isotopes explicitly.
@@
We further note that, from silicon core burning onward, the energy output in the core is not computed using each individual reaction but rather a quasi-statistical equilibrium (QSE) state between two elementary groups, around ${}^{28}\rm Si$ and ${}^{56}\rm Ni$. This is done by only taking into account the energy produced by reactions involving isotopes with $Z<14$, lower than ${}^{28}\mathrm{Si}$, and the ${}^{44}\rm Ti(\alpha,\gamma)^{48}\rm Cr$ reaction, which serves as the link between the 2 QSE groups of ${}^{28}\rm Si$ and ${}^{56}\rm Ni$ (for further detail see Hix et al. 1998; Hirschi et al. 2004). This approximation is applied due to the lack of stability of the heyney solver during silicon core burning of the models when using the energy output of individual reactions. Coupling the entire chemical composition to the structure equations as done by MESA (Paxton et al. 2011, 2013, 2018, 2019) could likely create a more stable environment. Also, correctly including the feedback on the temperature within the nuclear network would help to evolve the energy generation coupled with the network. However, both of these points go beyond the scope of this work. The energy losses by the electron captures, carried away by neutrinos, are included explicitly on top of the QSE calculation, thus allowing for a consistent contraction of the core in the final phases, where electron captures are the dominant energy sink in the core for the stellar masses of interest.

### v2

@@ 2.3 Impact on thermodynamic derivatives
Figure 3: Relative difference of the thermodynamic derivative $\nabla_{\rm ad}$ between the two EoSs, D-EoS and T-EOS, in log scale. The composition used is that of pure $\rm{}^{16}O$, $\overline{A}=16$ and $\overline{Z}=8$. The numbered zones represented are the same as in Figs. 1 and  2.
Between different EoSs the thermodynamic derivatives can change significantly. Specifically, we focus on the derivative needed to compute the adiabatic gradient:
@@
This must be obtained with sufficient accuracy, since $\nabla_{\rm ad}$ is a fundamental variable to determine whether a mass shell is unstable to convection.
@@ 3 Extension of the opacity coverage
The current implementation of opacities for the advanced phases in GENEC does not cover the full $(\rho,T)$ plane needed to evolve massive star models until core collapse. We shall use analytic extensions to estimate the opacities in these high-energy and relativistic regimes.
@@
In the following we detail the extensions employed to compute both $\kappa_{\rm cond}$ and $\kappa_{\rm rad}$ in the high-density and high-temperature regime, which is reached as stellar models evolve to core-collapse.
@@ 3.1 Radiative opacity
For most of the $(\rho,T)$ plane the radiative opacities are evaluated from tabulated values (see Ekström et al. 2012, for further details), which take into account many physical processes such as bound-bound transitions, bound-free transitions, free-free transitions and electron scattering to name just a few. The coverage of these tables generally only extend to temperatures of $10^{9}$ K. At higher temperatures, the opacity is dominated by electron scattering, but the Klein-Nishina corrections to Compton scattering lower it (Weaver et al. 1978), and at even higher temperatures, the effects of electron-positron pairs must be included in the opacity (Woosley et al. 2002). In the high-density regime, electron conduction is an important contribution to the opacity (see Itoh et al. 1985), and also electron degeneracy tends to reduce the opacity due to occupation of the electron phase space. To extend the computation of the radiative opacity to these regions, we apply the work of Poutanen (2017) to estimate the Rosseland mean opacity for electron scattering in the high-temperature regime. This opacity is expressed as
$\kappa_{\rm es}=\frac{\sigma_{T}N_{e}}{\rho}\frac{1}{\Lambda},$ (9)
where the parameter $\Lambda(T,\psi)$ is obtained from a fit in temperature and degeneracy parameter. The fit employed is,
@@ 3.2 Conductive opacity
The conductive opacities due to electrons are computed using the formulation of Iben (1975), given in the Appendix of that paper. These were already employed in the previous version of GENEC (Ekström et al. 2012); however, here the expressions have been modified to apply also in the final phases of stellar evolution. For this reason, we detail next the procedure used to compute $\kappa_{\rm cond}$. There are two main cases, the first being when the electrons are non-relativistic, roughly for $\rho<10^{6}\,\text{g\,cm}^{-3}$. In that case, a fit is developed based on the results of Hubbard & Lampe (1969), where they compute the thermal electron conductivity for mixtures of pure hydrogen, helium, and carbon. Thus, for densities smaller than $10^{6}\,\text{g\,cm}^{-3}$ the conductive opacity is written as
@@
$Z_{C}=\frac{1}{3}\sum_{i}Z_{i}^{2}X_{i}/A_{i}$ (16)
for all species beyond helium, ($T_{6}=T/10^{6}$). [^4] In the previous versions of GENEC, only a selection of all species above helium was included in $Z_{C}$. In particular, most of the species with $A\geq 56$, were not included. In the upgraded version, all species consistent with the nuclear reaction network used are for the computation of $Z_{C}$. The same comment applies to the quantities $Z_{\alpha}$ and $Z_{\beta}$ needed for the evaluation of $\kappa_{\rm cond}$ in Eq. (17). Here $\theta_{X}$, $\theta_{Y}$, $\theta_{C}$, and $f$ are fit parameters, dependent on thermodynamic quantities, as detailed in Iben (1975) to match the tabulated results of Hubbard & Lampe (1969). We note that the factor 1/3 in Eq. 16 is there to normalise the weight of $X_{\rm{}^{12}C}$ to 1 ($36^{2}/12=$ 3) because the computation for $\kappa_{\rm cond}$ in Hubbard & Lampe (1969) is only done for pure Carbon.
In the case of relativistic degenerate electrons, Iben (1975) proposes a fit for the results of Canuto (1970), which extend up to densities of $\rho\approx 10^{12}\,\text{g\,cm}^{-3}$. The approximation is
$\kappa_{\rm cond}=6.35\times 10^{-8}\times\frac{T_{6}^{2}Z_{\beta}}{\epsilon_{F}^{2.5}G}\times\frac{1}{(1+\epsilon_{F})}\ \mathrm{cm^{2}\,g}^{-1},$ (17)
where
$\epsilon_{F}$ $=\left[1+\left(\frac{\rho_{6}}{\mu_{e}}\right)^{2/3}\right]^{1/2}-1,$
$Z_{\beta}$ $=\sum_{i}Z_{i}^{2}X_{i}/A_{i},$
$\log G$ $=[(0.873-0.298\log Z_{\alpha}),$
$+(0.333-0.168\log Z_{\beta})M]\times\left[1-(1+\gamma)^{-0.85}\right],$
$Z_{\alpha}$ $=\sum_{i}Z_{i}^{2}X_{i}/A_{i}^{1/3},$
@@
$M$ $=\text{min}(1,0.5+\log\epsilon_{F}).$
[…]
3.3 Impact of updated opacity
To demonstrate the impact of the improved opacity calculation on the evolution of massive stars, we have computed a test model of initial mass $15M_{\odot}$ and solar metallicity, which is exactly the same as the model we present later in Sect. 5, except that the improved opacities of the previous sections have not been used. The new opacity routine only impacts the computation in the very final stages of evolution, where the density and temperature are high enough for the corrections to make any effect. In Fig. 4, we show the structure of the inner core for both of these models post-silicon core burning, and thus close to collapse.
Figure 4: Thermodynamic variables plotted for $15M_{\odot}$, with and without the opacity corrections. The model is plotted during silicon shell burning some minutes before collapse. The solid line shows the model using the opacity upgrades for the radiative and conductive opacity as detailed in Sect. 3. The dotted line shows the model run with the opacity previously implemented in GENEC. We display the compactness at an entropy per baryon $s=4$, computed using the definition in Eq. 24.
The correction to the opacity is quite consequential for the final structure of the model, specifically, in the innermost regions of the star. That can be seen in Fig. 4, which displays the structure of the inner 2 solar masses for the $15M_{\odot}$ computed using GENEC with (solid lines) and without (dashed lines) the opacity corrections. The opacity changes by 2 to 3 orders of magnitude in the inner core. The most important contribution in this change is due to the inclusion of all species above helium (and not only a selection of them; see footnote 4) in the conductive opacity correction. Without the correction the GENEC model has a lower opacity and this leads to a more efficient cooling of the core. Therefore to rise the temperature to the same value the core must contract more. We see that the central density, shown in Fig. 4 is slightly higher for the non corrected model for the same central temperature. As the core has contracted further the entropy jump represented at $s=4$ is at a smaller Lagrangian mass coordinate. This leads to a more compact core as parametrised by the compactness, $\xi_{s=4}$ [^5] See Sect. 5 for more discussion on this parameter. Eq. 24. While the original opacity yields $\xi_{s=4}=0.55$, one obtains $\xi_{s=4}=0.41$ for the model with corrected opacities. The applied corrections are thus crucial to determine the most accurate progenitor structure at collapse.
@@ 4 Implementation of an effective network
To reach the brink of core collapse, massive stars progressively neutronise their cores, reducing the electron fraction, $Y_{e}$, and with it the supporting degenerate electron pressure. Any reaction network accounting for the most advanced burning phases, must include (at least, approximately) the weak reactions that reduce the core electron fraction (mostly, electron capture onto elements of the iron group). In the previous work by Hirschi et al. (2004), GENEC models were computed post-silicon core-burning and even post-silicon shell burning, thus bringing them very close to collapse models. However, the network employed did not account for any decrease in $Y_{e}$. To extend the use of GENEC models as supernova progenitors, we have implemented electron capture in a reduced network to model the $Y_{e}$ decrease seen in stellar cores prior to collapse. Electron capture is especially important from silicon core burning to collapse (Woosley et al. 2002). Ideally, one would include a large set of reactions and isotopes to resolve with precision the $Y_{e}$ evolution. For instance, Rauscher et al. (2002) have employed a network of up to 2 20022002\,2002 200 isotopes, Farmer et al. (2016) using MESA computed the pre-supernova structure of models with various nuclear reaction networks of different sizes and Limongi & Chieffi (2018) used a network of roughly 300 isotopes.
In this work, we use a simple extension of the pre-existing $\alpha$-chain backbone network of GENEC. We approximate the effects of weak reactions that one would get in a larger network by taking one simple electron capture chain from $\rm{}^{56}Ni$ to $\rm{}^{56}Cr$. In this sense we create a similar network to the 21 species, detailed for example in Farmer et al. (2016), commonly refereed to in MESA literature (Paxton et al. 2011, 2013, 2015, 2018, 2019) as approx21_plus_co56. The exact network we use contains 25 species [^6] In the main sequence and helium core burning a larger network is used to capture carbon nitrogen oxygen, CNO, cycles correctly (see Ekström et al. (2012) for details). From carbon burning onwards, we use the 25-species network described here. and is presented in Fig. 5. Hereafter, we refer to this reduced network as GeValNet25. Compared with the approx21_plus_co56 network, we additionally include explicitly the iron group elements $\rm{}^{53}Fe$, $\rm{}^{55}Fe$, $\rm{}^{55}Co$, and $\rm{}^{57}Co$ to model the core neutronisation within the framework of GENEC.
@@
It should be noted that the last electron capture in this chain contains a double step, first from $\rm{}^{56}Fe$ to $\rm{}^{56}Mn$, and then to $\rm{}^{56}Cr$. This electron chain mimics the effect of a much larger network containing many electron reactions (cf. Aufderheide et al. 1994; Langanke & Martínez-Pinedo 2000). The rates used here were calibrated in a way that we reach similar core electron fractions as previous studies (Weaver et al. 1978; Woosley et al. 2002; Heger & Woosley 2010; Chieffi et al. 1998; Farmer et al. 2016; Paxton et al. 2018). It should be noted that direct beta reactions are not explicitly included in our calculations (differently from the inverse beta reactions, i.e. electron capture). Hence, there is no true balance between weak processes, implying that once the chain of electron captures begin, it unavoidably brings the core towards a composition dominated by the last isotope of the chain. In other words, should the core not collapse, in a finite time all the core would be composed of $\rm{}^{56}Cr$. Thus, the electron capture rates have been suitably adapted to become effective weak rates that account for the state close to approximate beta equilibrium in the core. [^7] Certainly, this is not a detailed beta equilibrium where every electron capture is balanced by the corresponding beta-decay, but rather, a global one, where the large values of the reaction rates allow that the integral of all beta decays approximately balances with the integral of all capture rates, facilitating an effective equilibrium where $dY_{e}/dt$ is very small.Admittedly, this approach introduces a simplification in modelling weak interaction physics, which facilitates the integration of these processes within the GENEC framework. Future updates could consider incorporating a selection of direct beta-decay reactions to approximate a global beta equilibrium. However, the sparse nature of the reaction network introduces uncertainties regarding the optimal sample of beta decays needed to accurately represent the underlying weak interaction processes. Energy changes resulting from effective electron capture are accounted for by applying the $Q$-value for these reactions, with an adjustment for the energy lost through neutrinos during weak interactions. The energy of the neutrinos is taken from the values found in Langanke & Martínez-Pinedo (2001). As pointed out in Heger et al. (2001b), due to the quasi beta equilibrium in the core, the values of the beta decay rates are immaterial, as long as they facilitate the effective equilibrium (i.e. as long as they are sufficiently large). Even variations by factors 2 to 5 in the beta-decay rates do not significantly change the nuclear evolution. We note that the choice of using an electron capture chain along $A=56$ until $\rm{}^{56}Cr$ is not by chance. In studies with larger networks and more species, the final $Y_{e}$ value commonly falls around $0.43-0.45$ (Woosley et al. 2002). If the simplified electron capture chain stopped at $\rm{}^{56}_{26}Fe$, then the lowest $Y_{e}$ attainable would be $Y_{e}^{\rm min}=26/56\approx 0.46$ (when all the matter in a core mass shell would be converted to $\rm{}^{56}_{26}Fe$), too high to match the results of larger and more complete networks. However, including $\rm{}^{56}_{24}Cr$ in the chain allows one to evolve to $Y_{e}^{\rm min}=24/56\approx 0.42$. Thus, a core containing a mixture of $\rm{}^{56}Fe$ and $\rm{}^{56}Cr$ can reach the electron fraction values generally found with sophisticated studies, even without explicitly accounting for the balancing effect of beta-decays. We cautiously point out that, with a reduced network, we may not provide a complete forecast of the nuclear composition at the brink of collapse.
@@
Figure 5: Network used in all GENEC models from core C-burning onwards. Squares filled with a plus sign denote isotopes explicitly evolved in the nuclear reaction network. Squares filled with a blue dot are isotopes implicitly included in the $\alpha$-chain. By this we mean that along the $\alpha$-chain are included not only the ($\alpha$, $\gamma$) reaction but also the ($\alpha$, $p$)($p$, $\gamma$) reactions without explicitly evolving the intermediate isotope abundance. With such a network, we get a good estimation of the energy generation rate (see Timmes et al. 2000, for details). Additionally, compared to the previous GENEC network, we have the approximate electron capture from $\rm{}^{56}Ni$ to $\rm{}^{56}Cr$ to mimic pre-supernova $Y_{e}$ evolution. This network is very similar to that of the 22-isotope network of Farmer et al. (2016); we note that in Fig. 1 of that paper only 21 isotopes are flagged in the 22-isotope network, but $\rm{}^{56}Co$ is added to the network making 22 in total. Compared with that network, we track explicitly extra iron group isotopes, namely $\rm{}^{57}Co$ along with $\rm{}^{53}Fe$ and $\rm{}^{55}Fe$, and thus in total we track 25 isotopes explicitly.
@@
We further note that, from silicon core burning onwards, the energy output in the core is not computed using each individual reaction but rather a quasi-statistical equilibrium (QSE) state between two elementary groups, around ${}^{28}\rm Si$ and ${}^{56}\rm Ni$. This is done by only taking into account the energy produced by reactions involving isotopes with $Z<14$, lower than ${}^{28}\mathrm{Si}$, and the ${}^{44}\rm Ti(\alpha,\gamma)^{48}\rm Cr$ reaction, which serves as the link between the 2 QSE groups of ${}^{28}\rm Si$ and ${}^{56}\rm Ni$ (for further detail see Hix et al. 1998; Hirschi et al. 2004). This approximation is applied due to the lack of stability of the heyney solver during silicon core burning of the models when using the energy output of individual reactions. Coupling the entire chemical composition to the structure equations as done by MESA (Paxton et al. 2011, 2013, 2018, 2019) could likely create a more stable environment. Also, correctly including the feedback on the temperature within the nuclear network would help to evolve the energy generation coupled with the network. However, both of these points go beyond the scope of this work. The energy losses by the electron capture, carried away by neutrinos, are included explicitly on top of the QSE calculation, thus allowing for a consistent contraction of the core in the final phases, where electron capture is the dominant energy sink in the core for the stellar masses of interest.